My father is four times as old as me. In 20 years, he will only be twice as old as me. How old is my father and how old am I?
This is what I have worked out so far but I am confused:
f=father's age and m=my age
f=4m
(f+20)=2(m+20)
(4m+20)=2(f+20)
4m+20=2f+40

Answer by josgarithmetic(39620)

You made a computation mistake.

f=4m, and f+20=2(m+20). From second equation,
f+20=2m+40
substitute for f,
4m+20=2m+40
4m=2m+20
2m=20
Answer: m=10 therefore f=40.